# Microsoft Research Montréal, Canada

### v1

@@ 1 Introduction
Robust access to trustworthy information is a critical need for society including implications for knowledge production, public health education, and promoting informed citizenry in democratic societies. Generative AI technologies such as large language models (LLMs) may enable new ways to access information and improve effectiveness of existing information retrieval (IR) systems. However, the long-term social implications of deploying these technologies in the context of information access are not yet well-understood. Existing research has focused on how these models may generate biased and harmful content [22, 11, 67, 151, 121, 227, 78] as well as the environmental costs [22, 159, 30, 232, 60, 160] of developing and deploying these models at scale. In the context of information access, Shah and Bender [178] have argued that certain framings of LLMs as “search engines” lack the necessary theoretical underpinnings and may constitute as a category error. In this current work, we present a broader perspective on the sociotechnical implications of generative AI for information access. Our perspective is informed by existing literature and aims to provide a summary of known challenges viewed through a systemic lens that we hope will serve as a useful resource for future critical research in this area. We present a summary of these implications next followed by recommendations for evaluation and mitigation later in this chapter.
@@ 2 Implications of generative AI for information access
Table 1: Overview of the consequences for information access from generative AI, the related mechanisms introduced by these AI technologies, and corresponding risks.
@@ 2.1 Consequences and mechanisms
In the context of information access, we identify five consequences of moral and social import of generative AI, and corresponding mechanisms, that we discuss next.
@@ 2.1.1 Consequence: Information ecosystem disruption
To reflect on the implications of generative AI on information access, we must consider the information ecosystem as a whole, and not constrain our discussion only to the application of these emerging technologies directly in IR systems. This ecosystem includes different actors and stakeholders such as information seekers, content producers, IR systems developers, advertisers, and other sociopolitical actors. While the information ecosystem is constantly evolving, generative AI holds the potential to significantly disrupt how each of these actors operate on their own and how they relate to other actors and stakeholders. This potential for disruption spans across how content is produced, consumed, monetized, and manipulated towards specific ends. By no means do we want to imply that these plausible changes are inherently bad but the scale of these potential disruptions across the ecosystem should motivate careful and thoughtful considerations before these technologies are deployed at scale. We discuss next some the underlying mechanisms introduced by generative AI that may contribute to these disruptions. We encourage the reader to view these mechanisms not just in isolation but to also consider how they may interact with each other and how that may impact the ecosystem over time.
@@
Mechanism: Content pollution Generative AI enables low-cost generation of derivative low-quality content at unprecedented scale. As a consequence, synthetic AI-generated content are rapidly flooding the web [102]. On Amazon, [^1] https://www.amazon.com/ these AI-generated content include scammy derivatives of existing publications [152, 118, 131] and fake travel guides [122]. On YouTube, [^2] https://www.youtube.com/ AI-generated videos are targeting children [4, 101, 119]. We are also witnessing a proliferation of news websites on the web that are almost entirely generated by AI [174] which are being surfaced in search results [50] and being funded by online ads [32]. Even reputable publishers have been reported to have published AI-generated articles under fake AI-generated author profiles [61]. Beyond news, other synthetic content such as AI-generated images are starting to pollute search results [13, 62]. According to another recent study [204], a “shocking” amount of content on the web today are machine translated text. Concerns have also been raised about LLMs serving as “Misinformation Superspreaders” [31, 155] as they make it trivially easy to inundate the web with politcally-motivated firehose of falsehood. [^3] https://en.wikipedia.org/wiki/Firehose_of_falsehood Hoel [101] points out that this AI pollution of our information ecosystems is a “tragedy of the commons” [96].
@@
Pollution of our information ecosystem at such scale has critical implications for people and society. When authoring a document requires significant time and effort then the quality, style, and comprehensiveness are factors that readers may consider to decide whether and how much to trust its content. However, when the cost of writing an extensive article approaches zero, it become significantly harder for the reader to make that decision. Furthermore, the increasing adoption of these same AI authoring tools by reputable publishers and content producers may homogenize the language and style of content on the web making it further difficult for readers to distinguish them from low-quality AI-generated content whose sole intent is to generate ad-revenue or to disinform. This pollution of the web is also a concern for future AI models that require large web-scale datasets to train on. Including AI-generated in the training data for new AI models may have significant negative impact on model performance, what has been referred to in the current literature as “Model collapse” [183, 136], “Model Autophagy Disorder” [14], and “Habsburg AI”. [^4] https://twitter.com/jathansadowski/status/1625245803211272194
@@
Mechanism: The “Game of telephone” effect LLMs have recently been employed in conversational search interfaces. In systems such as Bing Copilot, the LLM has access to relevant web search results that it can draw information from to generate appropriate responses to the user’s expressed information needs. In this scenario, the LLM performs a complex summarization task extracting relevant information from the retrieved documents to answer the search query. In doing so, the LLM now inserts itself between the user and the retrieved web results. This shifts the responsibility of inspecting the information in the documents and assessing their relevance, trustworthiness, and surrounding context from the user to the LLM. Factual errors and inconsistencies may arise between what the LLM produces and what is in the retrieved documents. Seeing the model through an anthropomorphic lens, these errors are sometimes referred to as “hallucinations”. A more technical view may see this as a noisy translation akin to the children’s game of telephone. [^5] https://en.wikipedia.org/wiki/Game_of_telephone These errors, often subtle and hard to spot, however may contribute to misinformation and reduce robustness of the information access system. While the LLM-generated responses may cite relevant documents, it is unlikely that users diligently click these links and verify the information in the response is indeed supported by said sources. Even if the LLM reproduces exact pieces of text from the source documents without error, taking these out of the context of the document may lead to unexpected negative consequences. Such examples have previously been reported [207] in context of extracted answers that search engines display on the search result pages (SERPs) as response to the user query. These issues may become more prevalent if conversational search interfaces become a popular way to access online information.
@@
In a more radical proposal, Metzler et al. [140] have suggested that LLMs could directly replace retrieval systems and respond directly to the user based on information in its training data. LLMs are trained to produce statistically plausible text sequences and any semblance to an information retrieval system is likely an important miscategorization of these models that we should be wary of [178]. The game of telephone effect is likely to be more intense when LLMs are expected to produce information from their training data and not just the in-context information in its input.
@@
This interjection of the LLM between the user and the search results may have other long term effects. These interfaces may alienate users from the practice of verifying information sources and make them less skilled over time at discerning online misinformation. If users get accustomed to information being presented neatly summarized disconnected from their sources, the critical cognitive skills necessary to distinguish between trustworthy and untrustworthy information may atrophy.
@@
Mechanism: Search engine manipulation New applications of LLMs to the IR stack have exposed new attack vectors. Prompt inject attacks [87, 133, 132] that try to blur the line between instructions and data have garnered specific interest. In these types of attacks, website owners may inject what looks like instructions to the LLM. When such documents are retrieved and included in the input of the LLM as augmentation, the LLM may mistake the injected prompt in the document content and be vulnerable to manipulation.
@@
Recently, LLMs have also found application in relevance labeling for search [203]. It is not well understood yet whether this may make the search engine vulnerable to improper ranking manipulation by website owners and search engine optimization experts. For example, one may employ the same, or similar, LLMs to reproduce the labeling scheme externally and then overfit their website content and design to achieve undue high predicted relevance against queries to rank higher on SERPs.
@@
Other attack vectors may include using LLMs to create effective content farms to manipulate ranking of web results or even use LLMs to artificially simulate users interacting with the search system to fake clicks and other user behavior signals that search engines depend on.
@@
Mechanism: Degrading retrieval quality LLMs may negatively impact search result quality not just by contributing to new attack vectors but more worryingly in some cases the negative effect may be a result of the LLM behaving exactly as it is supposed to. For example, one potential consequence of conversational search interfaces becoming significantly more prevalent than the classic SERP is that the quality of feedback from user behavior signals on SERP may significantly degrade. Historically, users of commercial web search engines have given the systems noisy implicit feedback through clicks and other actions on SERP. These clicks are one of the key secret sauce of any modern search systems. However, conversational interfaces may discourage direct user clicks on web results and at best provide much weaker satisfaction signal that may be gleaned from the users’ next utterance in the conversation. This over time may negatively impact the underlying retrieval quality making the systems worse.
@@
Mechanism: Direct model access Another important consideration is the implications of open foundation models [112]. While centralized systems have their own negative implications, as we discuss in §2.1.2, open access generative AI models without any access moderation also pose certain challenges. For example, there are many classes of harmful intents that systems should refuse to respond to for minimizing harm. This may include search queries seeking information on methods to self-harm or cause harms to others. Publicly accessible LLMs trained on large web corpus may produce irresponsible responses to such queries in absence of content moderation. Even if the model is trained to not respond to certain classes of queries it is likely that there will be leakage. Such leakage may also happen in the context of traditional search systems. However, in the latter case, all queries are typically instrumented which allows for post-hoc analysis and identification of critical gaps in the moderation system. Unfortunately, no such mitigation is possible once these generative AI models are released in the wild.
@@
Mechanism: The paradox of reuse Content producers and information access technologies share a critical inter-dependant relationship [138, 217]. Websites such as Wikipedia, [^6] https://www.wikipedia.org/ StackExchange, [^7] https://stackexchange.com/ and Reddit [^8] https://www.reddit.com/ produce critical content that are surfaced by information access platforms (e.g., web search engines) and contribute to making these platforms significantly more useful to their users. In return, these platforms have historically sent traffic back to these websites that contribute to their increased readership, subscription, and monetization. However, when these platforms stop directing traffic back to these websites—e.g., by instead surfacing the relevant content directly on the search result pages (SERPs)—the relationship becomes less symbiotic towards the content producers, a phenomenon Taraborelli [198] termed the “paradox of reuse”.
@@
The application of LLMs as conversational information access interfaces is likely to significantly intensify this problem. For example, LLMs such as ChatGPT [^9] https://chat.openai.com/ and Google Gemini [^10] https://gemini.google.com/app may gobble up large quantities of content from these websites as part of their training data and later regurgitate the same information without any attribution back to the source websites. Even when these models summarize information from multiple online sources with attribution—e.g., Bing Copliot [^11] https://www.bing.com/chat, they typically deemphasize the references and reduce the likelihood of the searcher clicking through to the source websites compared to the classic ten-blue-links interface. There is evidence [56] that support that this phenomenon is already happening at scale and is jeopardizing the “grand bargain at the heart of the web” [97].
@@ 2.1.2 Consequence: Concentration of power
Technology shapes and is shaped by the sociopolitical power structures within which it exists. Deliberation on the social consequences of any technology must therefore include critical consideration of how the technology, and general narratives about said technology, shifts power and re-architects and codifies structures of hierarchy and control. It must also include critical reflections on the politics and values of those in power to oversee what and how technology is built, especially when they reinforce hierarchy and authoritarianism (e.g. [76, 123, 63]). The 2024 edition of the World Economic Forum’s Global Risks Report [10] lists “technological power concentration” as one of the top global risks for the coming decade and as the biggest upward mover in their annual ranking of global risks compared to the previous year. A report [109] from the research institute AI Now [^12] https://ainowinstitute.org/ similarly asserts “the concentration of economic and political power in the hands of the tech industry—Big Tech in particular” as the core challenge posed by AI. They further note that not just the technologies but the narratives (both the hype and the fear-mongering) around them questionably bolsters claims of “foundational” advancements and their unassailable equivalence with scientific progress. These concerns are complemented by the discourses within the AI community, such as the observations by Birhane et al. [26] that the prominent values expressed and operationalized in top cited AI papers generally have implications in support of centralization of power. Even if platform owners acted accountably to civil society, the concentration of power and control in their hands makes them vulnerable to other actors, such as autocratic governments, and allows that power to be potential abused for oppressive and harmful intents.
@@
The popularization of generative AI is likely to further entrench Big Tech’s concentration of power as they emerge as some of the only institutions that have the resources to develop and deploy these technologies [114]. We must interrogate how the application of these technologies for information access may contribute to further concentration and growing inequities of wealth and power. Towards that goal, we next discuss three mechanisms in the context of generative AI that may contribute to concentration of power and control.
@@
Mechanism: Compute and data moat The development of generative AI is heavily reliant on the availability of large swaths of training data and large-scale computing power for training and deployment. Only a handful of institutions, largely in the private sector, own and control these necessary resources while simultaneously evangelizing AI as crucial geopolitical leverage and critical social infrastructure [109]. Increased access to these models have sometimes been touted as potential paths to mitigation [5, 187], where access may range from being heavily restricted over API to “fully open” models. The ability to download these models with their learned parameters allows others to further adapt for their own applications and open the door to more meaningful analysis and audit of these models. However, such “open access” also have severe limitations that we should recognize. The availability of the trained models does little to challenge Big Tech’s predominant vision of what AI fundamentally looks like, which would require dismantling the data and compute moat itself and turning them into public infrastructure. The availability of such infrastructure would allow a broader section of society to participate in the reimagination and development of diverse approaches to AI and not merely being forced to be satisfied with critiquing and finetuning artefacts produced by these institutions. This also illustrates the importance of existing institutions such as archives, libraries and universities that have reliable, historical data. Similarly, while the research community would benefit from easier access to models for critical studies and audit, it is significantly limiting unless that access is also extended to the user-facing systems in which these technologies are deployed and corresponding instrumentation data that provides context on how these systems are used by people and potential consequences.
@@
Mechanism: AI persuasion There is an emerging recognition of the dangers of AI persuasion [35, 39, 157, 65], which Burtell and Woodside [35] define as “a process by which AI systems alter the beliefs of their users”. AI systems may persuade users by appealing to their reason and argument (“rational persuasion”) or by exploiting their cognitive biases and heuristics (“manipulation”) [65]. El-Sayed et al. [65] identify six mechanisms of generative AI persuasion—namely (i) trust and rapport, (ii) anthropomorphism, (iii) personalisation, (iv) deception and lack of transparency, (v) manipulative strategies, and (vi) alteration of choice environment —and corresponding model features that contribute to these mechanisms. In the context of information access and advertising, these capabilities of generative AI can be powerful tools to hyper-target users and manipulate their behaviors.
@@
Modern online information access and communication platforms monetized with targeted advertising have ushered in an age of surveillance capitalism [238, 239]. These information access systems increasingly collect detailed user behavior data that allow them to build accurate user profiles for audience targeting. There is strong evidence that people are more likely to consume information that oppose their own personal views and beliefs when the latter employs language similar to their own political leaning [234]. So, combining this treasure trove of users’ private preferences and behavioral data with the capabilities of generative AI to produce persuasive language could create terrifyingly effective tools for mass behavioral manipulation. The impact of such pervasive algorithmic nudging [144] may be further pronounced over longer time periods from continuous interactions between the user and the system. Putting these capabilities in the hands of online platform owners, which typically tends to be large multinational for-profit institutions with largely hierarchical non-democratic internal governance structures, poses serious risks to functioning of democratic societies. All platforms moderate content posted or accessible through the platform [81] and in doing so they unavoidably impose their values on their users. But even without an explicit politic, these platforms by design try to affect consumer choices as part of their advertising business models. So, there is a direct incentive for these platforms to use generative AI to produce hyper-targeted highly-personalized persuasive advertisements to convince users to make certain buying decisions. When these platforms optimize for increasing user engagement out of their own commercial interests they may knowingly or unknowingly incentivize these generative AI models to be more persuasively rage-baiting [105].
@@
Mechanism: AI alignment To prevent generative AI models from producing harmful and offensive content recent research have focused on how to align these model outputs with “human values” [113, 195, 172, 70, 71]. Approaches such as reinforcement learning from human feedback (RLHF) [42, 237] have been effective in limiting certain types of problematic content from being produced. However, this approach presupposes some notions of desirable values and puts the burden of determining and enforcing them on the shoulders of platform owners. Any notions of universal values that determines what type of content these models should generate—or, not generate [206]—is highly contested [168, 25, 108, 165, 175]. Placing these decisions in the exclusive domain of the platform owners, especially in the absence democratic and civil society oversight, further concentrates power to these platforms.

### v2

# Sociotechnical Implications of Generative Artificial Intelligence for Information Access

@@ 1 Introduction
Robust access to trustworthy information is a critical need for society including implications for knowledge production, public health education, and promoting informed citizenry in democratic societies. Generative AI technologies such as large language models (LLMs) may enable new ways to access information and improve effectiveness of existing information retrieval (IR) systems. More efficient basic task execution with the help of LLMs can also enable people to focus on the more challenging aspects of information retrieval related tasks and research. However, the long-term social implications of deploying these technologies in the context of information access are not yet well-understood. Existing research has focused on how these models may generate biased and harmful content [23, 11, 69, 158, 124, 236, 80] as well as the environmental costs [23, 166, 31, 241, 61, 167] of developing and deploying these models at scale. In the context of information access, Shah and Bender [187] have argued that certain framings of LLMs as “search engines” lack the necessary theoretical underpinnings and may constitute as a category error.
In this current work, we present a broader perspective on the sociotechnical implications of generative AI for information access. Our perspective is informed by existing literature and aims to provide a summary of known challenges viewed through a systemic lens that we hope will serve as a useful resource for future critical research in this area. We present a summary of these implications next followed by recommendations for evaluation and mitigation later in this chapter.
@@ 2 Implications of generative AI for information access
Table 1: Overview of potential negative consequences for information access from generative AI, the related mechanisms introduced by these AI technologies, and corresponding risks.
@@ 2.1 Consequences and mechanisms
In the context of information access, we identify five potential categories of negative consequences of generative AI, and corresponding mechanisms, that we discuss next.
@@ 2.1.1 Consequence: Information ecosystem disruption
To reflect on the implications of generative AI on information access, we must consider the information ecosystem as a whole, and not constrain our discussion only to the application of these emerging technologies directly in IR systems. This ecosystem includes different actors and stakeholders such as information seekers, content producers, IR systems developers, advertisers, and other sociopolitical actors. While the information ecosystem is constantly evolving, generative AI holds the potential to significantly disrupt how each of these actors operate on their own and how they relate to other actors and stakeholders. This potential for disruption spans across how content is produced, consumed, monetized, and used towards specific ends. By no means do we want to imply that these plausible changes are inherently bad but the scale of potential disruptions across the ecosystem should motivate careful and thoughtful considerations before these technologies are deployed at scale. We discuss next some the underlying mechanisms introduced by generative AI that may contribute to these disruptions. We encourage the reader to view these mechanisms not just in isolation but to also consider how they may interact with each other and how that may impact the ecosystem over time.
@@
Mechanism: Content pollution Generative AI enables low-cost generation of derivative low-quality content at unprecedented scale. As a consequence, synthetic AI-generated content is rapidly and very widely appearing on the web [104]. On Amazon, [^1] https://www.amazon.com/ AI-generated content includes scammy derivatives of existing publications [159, 121, 135] and fake travel guides [125]. On YouTube, [^2] https://www.youtube.com/ AI-generated video creators have targeted children [4, 103, 122]. We are also witnessing a proliferation of news websites almost entirely generated by AI [182], which are being surfaced in search results [51] and funded by online ads [33]. Even reputable publishers have reportedly published AI-generated articles under fake AI-generated author profiles [62]. Beyond news, other synthetic content such as AI-generated images is starting to pollute search results [13, 63]. According to another recent study [213], a “shocking” amount of content on the web today is machine-translated text. The promise of machine translation is that it could make more content accessible to wider audiences. However, it also amplifies the influence of (sometimes questionable-quality) language technology choices. For example, Thompson et al. [213] found that more low quality content—rather than high quality content—was machine translated into lower resource languages, likely with the goal of generating ad revenue. Concerns have also been raised about LLMs potentially serving as “Misinformation Superspreaders” [32, 162] as they make it trivially easy to inundate the web with “firehoses of falsehoods”. [^3] https://en.wikipedia.org/wiki/Firehose_of_falsehood Hoel [103] points out that AI pollution of our information ecosystems is a “tragedy of the commons” [98].
@@
Pollution of our information ecosystem at such scale has critical implications for people and society. When authoring a document requires significant time and effort then the quality, style, and comprehensiveness are factors that readers may consider in deciding whether and how much to trust its content. However, when the cost of writing an extensive article approaches zero, it becomes significantly harder for the reader to make that decision. They may not be able to distinguish between an article created based on extensive research, fact checking and thoughtful writing practices versus one generated instantly based on a short user prompt. Furthermore, the increasing adoption of these same AI authoring tools by reputable publishers and content producers may homogenize the language and style of content on the web, making it even more difficult for readers to distinguish them from low-quality AI-generated content whose sole intent is to attract ad revenue or to mislead. Such web pollution is also a concern for future AI models that require large web-scale datasets to train on. Including AI-generated content in the training data for new AI models may have significant negative impact on model performance, what has been referred to as “Model collapse” [192, 142], “Model Autophagy Disorder” [14], and “Habsburg AI”. [^4] https://twitter.com/jathansadowski/status/1625245803211272194
@@
Mechanism: The “Game of telephone” effect LLMs have recently been employed in conversational search interfaces. In systems such as Bing Copilot, the LLM has access to relevant web search results from which it can draw information to produce appropriate responses for the information needs expressed by a user. In this scenario, the LLM performs a complex summarization task extracting relevant information from the retrieved documents to answer the search query. In doing so, the LLM now inserts itself between the user and the retrieved web results. This shifts the responsibility of inspecting the information in the documents and assessing their relevance, trustworthiness, and surrounding context from the user to the LLM. Further, factual errors and inconsistencies may arise between what the LLM produces and what is in the retrieved documents. Seeing the model through an anthropomorphic lens, these errors are sometimes referred to as “hallucinations”. A more technical view may see this as a noisy translation akin to the children’s game of telephone. [^5] https://en.wikipedia.org/wiki/Game_of_telephone Such errors, often subtle and hard to spot, may contribute to misinformation and reduce robustness of the information access system. While the LLM-generated responses may cite relevant documents, it is unlikely that users diligently click the provided links and verify the information in the response is indeed supported by said sources. Even if the LLM reproduces exact pieces of text from the source documents without error, taking these out of the context of the document may lead to unexpected negative consequences. Such examples have previously been reported [216] in context of extracted answers that search engines display on the search result pages (SERPs) as response to the user query. These issues may become more prevalent if conversational search interfaces become a popular way to access online information.
@@
In a more radical proposal, Metzler et al. [146] have suggested that LLMs could directly replace retrieval systems and respond directly to the user based on information in their training data. LLMs are trained to produce statistically plausible text sequences and any semblance to an information retrieval system is likely an important mis-categorization of these models that we should be wary of [187]. The game of telephone effect is likely to be more intense when LLMs are expected to produce information from their training data and not just the in-context information in its input.
@@
The interjection of the LLM between the user and the search results may have other long term effects. These interfaces may disincentivize users from the practice of verifying information sources and make them less skilled over time at discerning online misinformation. If users get accustomed to information being presented neatly summarized and disconnected from original sources, the critical cognitive skills necessary to distinguish between trustworthy and untrustworthy information may atrophy.
@@
Mechanism: Search engine manipulation New applications of LLMs to the IR stack have exposed new attack vectors. Prompt injection attacks [89, 138, 137] that try to blur the line between instructions and data have garnered specific interest. In these types of attacks, website owners may inject what looks like instructions to the LLM. When such documents are retrieved and included in the input of the LLM as augmentation, the LLM may mistake the injected prompt in the document content and be vulnerable to manipulation.
@@
Recently, LLMs have also found application in relevance labeling for search [212]. It is not well understood yet whether this may make the search engine vulnerable to improper ranking manipulation by website owners and search engine optimization experts. For example, one may employ the same, or similar, LLMs to reproduce the labeling scheme externally and then adapt their website content and design to achieve undue high predicted relevance against queries to rank higher on SERPs.
@@
Other attack vectors may include using LLMs to create effective content farms at low cost to manipulate the ranking of web results, or even use LLMs to artificially simulate users interacting with the search system to fake clicks and other user behavior signals, such as reformulations, that search engines depend on.
@@
Mechanism: Degrading retrieval quality LLM usage can negatively impact search result quality in a number of (indirect) ways. LLMs can contribute to new attack vectors, but more worryingly, in some cases the negative effect may be a result of the LLM behaving exactly as it is supposed to. For example, one potential consequence of using conversational search interfaces, is that the quality of feedback from user behavior signals on SERPs may significantly degrade. Historically, users of commercial web search engines have given search systems noisy implicit feedback through clicks and other actions on SERPs. These actions are one of the key secret sauce of any modern search systems.
However, conversational interfaces may discourage direct user clicks on web results and at best provide much weaker satisfaction signal that may be gleaned from the users’ next utterance in the conversation. This over time may negatively impact the underlying retrieval quality. This makes it important to invest in methods that can infer user satisfaction with high certainty from the natural language conversations. However, methods for such signal interpretation are not yet at the level necessary to mitigate these impacts.
@@
Mechanism: Direct model access Another important consideration is the implications of open foundation models [115]. While centralized systems have their own negative implications, as discussed in §2.1.2, open access generative AI models without any access moderation also pose certain challenges. For example, there are many classes of harmful intents that systems should refuse to respond to. This may include search queries seeking information on methods to self-harm or cause harms to others, or requests to generate harmful (and sometimes illegal) content such as child sex abuse material (CSAM) or non-consensual intimate information (NCII). Publicly accessible LLMs trained on large web corpora may produce such irresponsible content in the absence of moderation. Even if a model is trained to not respond to certain classes of queries, it is likely that there will be leakage, and the safety alignment may also be compromised if the model is further finetuned [176]. Such leakage may also happen in the context of traditional search systems. However, in the latter case, all queries are typically logged, allowing for post-hoc analysis and identification of critical gaps in the moderation system. Unfortunately, no such mitigation is possible once these generative AI models are released into the wild.
@@
Mechanism: The paradox of reuse Content producers and information access technologies are critically inter-dependant [144, 226]. Websites such as Wikipedia, [^6] https://www.wikipedia.org/ StackExchange, [^7] https://stackexchange.com/ and Reddit [^8] https://www.reddit.com/ produce critical content that is surfaced by information access platforms (e.g., web search engines) and contribute to making these platforms significantly more useful to their users. In return, these platforms have historically sent traffic back to the websites that contributes to their increased readership, subscriptions, and monetization. However, when search platforms stop directing traffic back to websites—e.g., by instead surfacing relevant content directly on the search result pages (SERPs)—the relationship becomes less symbiotic towards the content producers, a phenomenon Taraborelli [207] termed the “paradox of reuse”.
@@
The application of LLMs as conversational information access interfaces is likely to significantly intensify this problem. For example, LLMs such as ChatGPT [^9] https://chat.openai.com/ and Google Gemini [^10] https://gemini.google.com/app may gobble up large quantities of content from websites as part of their training data and later regurgitate the same information without any attribution back to the sources. Even when models summarize information from multiple online sources with attribution—e.g., Bing Copliot [^11] https://www.bing.com/chat, they typically de-emphasize the references and reduce the likelihood of the searcher clicking through to the source websites as compared to the classic ten-blue-links interface. There is evidence [57] to suggest that this phenomenon is already happening at scale and is jeopardizing the “grand bargain at the heart of the web” [99].
@@ 2.1.2 Consequence: Concentration of power
Technology shapes and is shaped by the sociopolitical power structures within which it exists. The 2024 edition of the World Economic Forum’s Global Risks Report [10] lists “technological power concentration” as one of the top global risks for the coming decade and as the biggest upward mover in their annual ranking of global risks compared to the previous year. Deliberation on the social consequences of any technology must therefore include critical consideration of how the technology, and general narratives about said technology, shifts power and re-architects and codifies structures of hierarchy and control. In this context, the politics and values of those in power to oversee what and how technology is built or regulated, especially when they reinforce hierarchy and authoritarianism (e.g. [78, 126, 64]), becomes important to consider.
A report [112] from the research institute AI Now [^12] https://ainowinstitute.org/ similarly asserts “the concentration of economic and political power in the hands of the tech industry—Big Tech in particular” as the core challenge posed by AI. They further note that not just the technologies but the narratives (both the hype and the fear-mongering) around them questionably bolster claims of “foundational” advancements and their unassailable equivalence with scientific progress. These concerns are complemented by the discourses within the AI community, such as the observations by Birhane et al. [27] that the prominent values expressed and operationalized in top cited AI papers generally have implications in support of centralization of power. Even if platform owners act accountably to civil society, the concentration of power and control in their hands makes them vulnerable to other actors, such as autocratic governments, and allows that power to be potentially abused for oppressive and harmful intents.
@@
The popularization of generative AI can concentrate that power within large companies, since they emerge as some of the only institutions with the resources to develop and deploy these technologies [117]. The application of these technologies for information access may contribute to further concentration and growing inequities of wealth and power; we discuss three mechanisms in the context of generative AI that may contribute to concentration of power and control.
@@
Mechanism: Compute and data moat The development of generative AI is heavily reliant on the availability of large swaths of training data and large-scale computing power for training and deployment. Only a handful of institutions, largely in the private sector, own and control these necessary resources while simultaneously evangelizing AI as crucial geopolitical leverage and critical social infrastructure [112]. Increased access to these models has sometimes been touted as potential paths to mitigation [5, 196], where access may range from being heavily restricted over API to “open weight” models [134]. The ability to download models with their learned parameters allows others to further adapt for their own applications and opens the door to more meaningful analysis and audit of these models. However, such “open access” also leads to severe limitations that we should recognize. The availability of the trained models does little to challenge the predominant visions put forth by large technology companies of what AI fundamentally should look like.
One potential direction would be to dismantle the data and compute moat by turning them over from private ownership into public infrastructure for independent researchers and developers and those affiliated with smaller institutions. This also illustrates the importance of existing institutions such as archives, libraries and universities that have reliable, historical data. The availability of public computer infrastructure would allow a broader set of developers to participate in the reimagination and development of diverse approaches to AI and not merely being forced to be satisfied with critiquing and finetuning artefacts produced by other institutions. However, there is no guarantee that without careful planning and incentives, a proliferation of smaller projects will lead to transformative new or more sustainable results.
Democratizing the control over computational resources provides a mechanism of checks and balances on the future directions of AI systems, and may allow for challenges to popular narratives and expectations about generative AI such as exponential growth in model size over time. Infrastructure is however also bound to the particular governing system, and local underlying goals and processes. Larger investments in existing research institutes, or new alternative companies or non-profits might in certain cases lead to faster results.
Similarly, the research community would benefit from easier access to industry models and APIs for critical studies and auditing. However, access to models or APIs alone is significantly limiting unless that access is also extended to the user-facing systems in which these technologies are deployed. The corresponding instrumentation data would provide context on how these systems are used by people and potential consequences. This can lead to practical privacy and security questions for platform teams. Practical support for decision making and for example the creation of standards to de-risk those concerns can help alleviate some of those concerns.
@@
Mechanism: AI persuasion There is an emerging recognition of the dangers of AI persuasion [36, 40, 164, 67], which Burtell and Woodside [36] define as “a process by which AI systems alter the beliefs of their users”. AI systems may persuade users by appealing to their reason and argument, or by using their cognitive biases and heuristics [67]. El-Sayed et al. [67] identify six mechanisms of generative AI persuasion—namely (i) trust and rapport, (ii) anthropomorphism, (iii) personalization, (iv) deception and lack of transparency, (v) manipulative strategies, and (vi) alteration of choice environment —and corresponding model features that contribute to these mechanisms. In the context of information access and advertising, these capabilities of generative AI can be powerful tools to hyper-target users and steer their behaviors.
@@
Modern online information access and communication platforms monetized with targeted advertising have been said to usher in an age of surveillance capitalism [247, 248]. Information access systems increasingly collect detailed user behavior data that allow them to build accurate user profiles for audience targeting. There is strong evidence that people are more likely to consume information that opposes their own personal views and beliefs when the it employs language similar to their own political leanings [243]. So, combining users’ private preferences and behavioral data with the capabilities of generative AI to produce persuasive language could create worrying tools for mass behavioral manipulation. The impact of such pervasive algorithmic nudging [151] may be further pronounced over longer time periods from continuous interactions between the user and the system. Putting these capabilities in the hands of online platform owners, which typically tend to be large multinational for-profit institutions with largely hierarchical non-democratic internal governance structures, poses serious risks to functioning of democratic societies. At the same time, platforms must make decisions about what is acceptable on their platforms to avoid negative user experiences, spam, unwelcoming behavior, and other negative occurrences beyond those outlined in legal compliance alone. Platforms moderate content posted or accessible through the platform [83] and in doing so they unavoidably impose implementations of values on their users, or the values incentivized by, say, advertising needs or other business model related motivations. For ads, this may mean an incentive to use generative AI to produce hyper-targeted highly-personalized persuasive advertisements which convince users to make certain buying decisions. For content, when platforms optimize for increased user engagement, they may knowingly or unknowingly incentivize generative AI models to be producing highly charged content, such as “rage-bait” [107], because it tends to be more persuasive and engaging.
@@
Mechanism: AI alignment To prevent generative AI models from producing harmful and offensive content, recent research has focused on how to align model outputs with “human values” [116, 204, 180, 72, 73]. Approaches such as reinforcement learning from human feedback (RLHF) [43, 246] have been effective in limiting certain types of problematic content from being produced. However, this approach presupposes some notions of desirable values and puts the burden of determining and enforcing them on the shoulders of platform / model developers. Any notions of universal values that might determine what type of content these models should generate—or, not generate [215]—is highly contested [175, 26, 111, 172, 183]. Placing these decisions in the exclusive domain of the platform developers, especially in the absence of democratic and civil society oversight, further concentrates power and responsibility. This is not an argument against content moderation itself but against the centralization of control over it without civil oversight or broader societal participation. As a pragmatic example, platforms may not necessarily have the necessary knowledge in-house, making it imperative for them to make successful connections to outside expertise.